how do you get these answers?

23. (10 points) An electronics store sells tablets and laptops that are currently supplied by two different manufacturers in Asia. They are considering aggregating these orders by obtaining both products from a single supplier who will charge $10,000 per delivery. Annual holding costs are 25%. The monthly demand and unit cost for each product are summarized below. Tablets Laptops Monthly Demand 12,000 5000 Unit Cost of Product $100 $400 If you aggregate all products, how many orders should be placed per year? a. 21.91 # orders per year- b. With the order frequency found in part a, how many of each product would be irn each order? Tablets Laptops Number in Each Order c With the order freq fofd the folowing values
-a 1.91 # orders per year With the order frequency found in part a, how many of each product would be in each order? b. Tablets aptops Number in Bach Order 57 2.342338.48 With the order frequency found in part a, find the following values. c. Annual Ordering Cost $,0O Annual Inventory Cost for Tablets Anual Inventory Cost of Laptops $136 124 38 28. 25 Total Annual Cost-
Answer #1
Table Demand , Dt = Monthly demand * 12 =12*12000 144000 Demand, Laptop, DL = 5000*12 60000
Price $            100.00 Price $            400.00
Holding cost , Ht $              25.00 Holding cost , Hl $            100.00
(25% of price) (25% of price)
New Shipping cost, S = $      10,000.00
Part a) Optimal joint order frequency, N
n* = SQRT((DtHt + DlHl)/(2S*))
DtHt + DlHl = 144000*25+60000*1000 9600000
2S 20000
n*= sqrt (9600000/20000) 21.91
Part b) Order Quantity for Tablet , Qt = Dt/n* = 144000/21.91 6572.34 Order Quantity for B, Ql = Dl/n*= 60000/21.91 2738.48
Part c) Inventory cost for tablet = (Q/2*Ht) = (6572.34/2)*25 $      82,154.25 Inventory cost = (Ql*Hl)/2 =(2738.48/2*100) $      136,924.0
Total inventory cost current strategy = 82154.12+136924 $   219,078.25
Annual ordering cost = 21.91*10000 $   219,100.00
Total Cost = Total inventory cost + Annual ordering cost $   438,178.25